Circular No. 6347

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996M IN ANONYMOUS GALAXY R. H. McNaught, Anglo-Australian Observatory, reports his discovery

of a supernova (mag ∼ 18), found on a U.K. Schmidt telescope R plate taken on Mar. 19.7 UT by K. S. Russell and D. M. James. SN 1996M is located 21′′ due west of the center of the host galaxy (in a faint knot on the edge of the galaxy), at α = 12^{h}51^{m}17^{s}.73, δ = −13^{o}01′11′′.7 (equinox 2000.0). A nearby star of mag ∼ 17.5 has position end figures 20^{s}.57, 00′59′′.4. No image appears at the position of SN 1996M in several surveys, the latest being in 1994 June.

P. Garnavich, Harvard-Smithsonian Center for Astrophysics, reports that an image obtained by N. Caldwell at the Mt. Hopkins 1.2-m telescope on Mar. 21 UT shows the supernova at V ∼ 18.2.

S. Benetti, European Southern Observatory (ESO); and N. Caon, Space Telescope Science Institute, report: “Inspection of a preliminarilyreduced CCD spectrum (range 375.7–695.2 nm, resolution 1.5 nm), obtained on Mar. 22.19 UT with ESO 3.6-m telescope (+ EFOSC1), confirms this object as a type-II supernova, ∼ 50 days after maximum. Strong Hβ and Hα lines with P-Cyg profiles are superimposed on a red continuum. Ca II (H and K), Ti II 455.0-nm, Fe II 501.8- and 516.9-nm, and Na I D lines are also present with P-Cyg profiles. The expansion velocity deduced from the Hβ, Hα, and Na I D absorption is ∼ 7000 km/s. The Fe II lines show a slower expansion velocity (∼ 4500 km/s). The recession velocity of the parent galaxy (measured from emission lines in the galaxy spectrum) is 6000 km/s. Preliminary photometry: V = 18.1, V − R = +0.5.”

L. Wang and J. C. Wheeler, University of Texas, write: “We have observed SN 1996M at the McDonald Observatory 2.1-m telescope (+ imaging grism instrument). Preliminary inspection of the spectrum (range 425.4– 814.2 nm, resolution 2.5 nm) confirms this object as a type II supernova close to maximum light. The Hα and Hβ lines are strong and show P-Cyg profiles. The FWZI of the Hα line shows the fastest expansion velocity of the ejecta to be ∼ 10 000 km/s. Other lines detected are He I/Na I at 587.6 nm and Fe II at 511.0 nm.”

SUPERNOVA 1996L IN ESO 266-G10 Corrigendum. On IAUC 6346, line 3, for J. M. James. read D. M.

James.

1996 March 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green